OpenStudy (anonymous):

One angle of an equilateral triangle measures (4x-20)degrees. What is the value of x?

hartnn (hartnn):

Angle of equilateral triangle are all 60 so 4x-20=60 can u solve this to find x ?

OpenStudy (nubeer):

if equilateral triangle .. all angles will be same so 3(4x-20) =180 12x -60 = 180 12x = 240 x = 20

OpenStudy (anonymous):

Thanks, you guys are awesome, i suck at math.